OpenStudy (anonymous):

Please help me with this calculus problem! Differentiate sin^2(cos^3x+tanx)

OpenStudy (anonymous):

Any multiple choice option?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

I already had that open :P, it just didnt seem right...

OpenStudy (anonymous):

I think it is somewhat right.

OpenStudy (anonymous):

So, I should follow that?

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

sin^(2)(cos^3x+tanx) use chain rule g(x) = x^(2) g'(x) = 2x s(x) = sin(cos^3x+tanx) s'(x) = (cos^3x+tanx)' = -3(cos^(2)(x))sin(x) + sec^(2)(x) Now we have to differentiate both cos^(3)x and tan(x) l(x) = x^(3) l'(x) = 3x^(2) m(x) = cos(x) m'(x) = -sin(x) o(x) = tan(x) o'(x) = sec^(2)(x) Now we put it all together f'(x) = 2sin(cos^3x+tanx) * -3((cos^(2)(x))sin(x) + sec^(2)(x)) f'(x) = -6sin(cos^3x+tanx)(cos^(2)(x))sin(x) + sec^(2)(x))

OpenStudy (anonymous):

you can probably simplify it more but meh

OpenStudy (anonymous):

Have you clicked the link?

OpenStudy (anonymous):

note the derivative of cos^(3)x = -3(cos^(2)(x))sin(x)

OpenStudy (anonymous):

I hope you can follow what I did to solve this

OpenStudy (anonymous):

Yeah, dw about it

OpenStudy (anonymous):

ok then I wont :\